1. Terbutaline comes as a 1-mg/ml ampule. How much do you need to give a 0.25mg SQ dose?

2. Using a schedule for Albuterol syrup of 0.1 mg/kg and prepared strength mixture of 2mg/5ml, how much is needed for a 44lb child?

3. Epoprostenol (Veletri) is a selective pulmonary vasodilator that is nebulized via a mesh nebulizer. The pharmacy prepares standard concentrated syringes of 1.5mg in 50ml syringes. The schedule dose is 50mg/kg/min. The patients IBW is 140.5lbs

What is the concentration per ml of each syringe in mug/ml?

What is the patients weight in Kg?

What is the administration dose in mcg/min?

What is the administration dose in ml/hr?

How much time will it take to administer the entire 50ml syringe?
Answer #1
  1. Terbutaline available = 1mg/ml ampule.

For 0.25mg SQ dose required ml= 0.25 ml

  1. Order 0.1 mg/kg

Available strength = 2mg/5ml= 0.4 mg /ml

Wt of child= 44 lb= 44/2.2=20 kg

As per wt of child the dose will be = 0.1 X 20= 2 mg

For 2 mg required ml= 2/0.4=5 ml will be required dose

  1. Available= 1.5mg in 50ml syringes.

The schedule dose is 50mg/kg/min.

The patients wt= 140.5lbs= 63.86 kg

Dose as per wt of patient = 50 X 63.86 mg/min= 3193 mg/min

Concentration per ml of each syringe in mg/ml= 1.5/50= 0.3 mg /ml

Patient’s weight in Kg= 140.5/2.2= 63.86 kg

What is the administration dose in mcg/min?

The schedule dose is 50mg/kg/min.

The patients wt= 140.5lbs= 63.86 kg

            Dose as per wt of patient = 50 X 63.86 mg/min

                                                      = 3193 mg/min

What is the administration dose in ml/hr?

3193 mg in ml= 3193X 3.33=10643 ml/min= 177.38 ml/hr

How much time will it take to administer the entire 50ml syringe?

177.38 ml required time = 60 min

1 ml= 60/177.38

Hence for 50 ml required time = (60/177.38) X 50= 16.91 min= 17 min